A dart hits the circular dartboard shown below at a random point. Find the probability that the dart lands in the shaded square region. The radius of the dartboard is 4in, and each side of the shaded region is 2in.
Use the value 3.14 for pi. Round your answer to the nearest hundredth.
1
Expert's answer
2017-07-31T14:30:06-0400
Total area=Area of the bigger circle= Ac=Pi*(radius of the dashboard)^2=3.14*4*4=50.24 square in.
Desired area=Area of the shaded smaller square= As=2*2=4 square in.
Then Geometric Probability= Area of the shaded smaller square/Area of the bigger circle=As/Ac=4/50.24 is approximately equal to 0.08.